OpenStudy (anonymous):

Find the complete solution for 3x - 2 ≡ 7 (mod 11) and 5 ≡ 4x -1 (mod 9). @MIT 18.01 Single …

OpenStudy (anonymous):

3x = 9 (mod 11) x = = 3(mod 11) 4x = 6 (mod 9) x = 1 1/2 (mod 9)